OpenStudy (anonymous):

Janine made a cylindrical vase in which the sum of the lateral area and area of one base was about 3000 square centimeters. The vase had a height of 50 centimeters. Find the radius of the vase. Explain the method you would use to find the radius.

OpenStudy (thefurball):

\[Area _{Lateral} = 2\pi rh\]\[Area _{base} = \pi r^{2}\]Sum of those two equations:\[2\pi r h+\pi r^{2} = 3000\]in this case, h is 50, so \[100\pi r + \pi r^{2}=3000\]Can you solve for r from there?

OpenStudy (anonymous):

Would it be 324? @thefurball

OpenStudy (anonymous):

Yes as he said; Area Lateral =2Trh. Areaxbase, =tTr^2. 2Trh + Tr^2 = 3000. Lateral x base/ Steps however are: for fractions, simply divide, then use the answer in the next part. complete it in segments, first the square root, then add the other top of the fraction, then divide the whole thing by the bottom. btw you'll have to type it in twice, first using the - sign, then the + sign. this means you'll end up with two answers, so the logically ridiculous answer (ex: negative or too small) is rejected. if you get a negative within the square root however, there'll be an error. it doesn't necessarily mean that the answer is wrong, but you just won't be able to use this quadratic formula method.... but yes in conclusion your answer is 324. Hope this helps. Have a nice day, tag me on anymore you need help with.